OpenStudy (anonymous):

Explain how to find the sum -65 + (+45+using a number line,

OpenStudy (cggurumanjunath):

-65+(+45)=-65+45=-20

OpenStudy (anonymous):

I Put (45) I meant 41

OpenStudy (cggurumanjunath):

-65+(+41)=-65+41=-24

OpenStudy (anonymous):